# Divergence-Kernel method for linear responses and diffusion models

### v1

@@ 5. A simple forward-only diffusion model
where the true parameters $\gamma_{true}:=[\gamma_{true}^{0},\ldots,\gamma_{true}^{N_{\gamma}}]$. The samples are $y:=x_{T}$ obtained at the end of independent paths. Denote the empirical measure, which we try to match, as
$\begin{split}P(\cdot):=\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta_{y_{k}}(\cdot).\end{split}$
Our framework can work on more general data, but here we restrict ourselves to the simple case, so that after optimization, we can check our result with the true parameters.
We want to fit several unknown parameters $\gamma=[\gamma^{0},\ldots,\gamma^{N_{\gamma}}]$ in an SDE such that its marginal distribution matches the data. The prototype of the SDE is
@@
Denote the density at $T$ by $h^{\gamma}$ and measure by $H^{\gamma}$ – this is the distribution that we want to optimize. For each $\gamma$ during optimization, we simulate $L$-many paths by this SDE, and we label each path by $l$, in particular, the terminal location is denoted by $x_{T,l}$. In future research, we can seek to use more flexible prototypes, such as neural networks. However, note that the gradient used in the optimization is not computed by backpropagation; rather, it only involves a covector process that runs forward in time and costs much less memory. We also remark that we can choose to optimize the stationary distribution using ˜Remark, which may be more efficient for some cases.
The difference between the generated distribution and the data is measured by the KL-divergence
$\begin{split}D_{\text{KL}}(P\parallel H^{\gamma}):=\int_{y\in{\mathbb{R}^{M}}}\log{\frac{P(dy)}{H^{\gamma}(dy)}}\,P(dy).\end{split}$
Its parameter-derivative is
$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta H^{\gamma}(dy)}{H^{\gamma}(dy)}\,P(dy)=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta h^{\gamma}}{h^{\gamma}}(y)\,P(dy)\\ =-\int_{y\in{\mathbb{R}^{M}}}\delta\log h^{\gamma}(y)\,P(dy)=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta\log h^{\gamma}(y_{k}).\end{split}$
With the new formula in Section˜1.1, we have an expression of $\delta\log h^{\gamma}$ at $x_{T,l}$’s, which enables the gradient optimization on this KL divergence.
Then we need to decide how to compute $\delta\log h^{\gamma}$ at the prescribed locations $y_{k}$’s. It is difficult to obtain paths ending exactly at $y_{k}$, so we need an approximation. We approximate $y_{k}$ by its $N_{neighbor}$-nearest neighbors in $\{x_{T,l}\}_{l=1}^{L}$, and therefore $\delta\log h^{\gamma}(y_{k})$ is approximated by the average of the divergence-kernel expression in Section˜1.1 over those $N_{neighbor}$ paths. Hence,
$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})\approx\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}):=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\frac{1}{N_{neighbor}}\sum_{x_{T,l}\in N(y_{k})}\mathcal{T}(x_{T,l}),\end{split}$
where $N(y_{k})$ is the neighbors of $y_{k}$, and $\mathcal{T}(x_{T,l})$ is the expression in Section˜1.1 on the path ending at $x_{T,l}$. After approximately computing the parameter-derivative, we update the parameter by gradient descent,
$\begin{split}\gamma_{n+1}=\gamma_{n}-\eta\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}).\end{split}$
@@
5.1. A 1-d problem with 2 parameters
First, we try a 1-dimensional problem. The SDE is
@@ 5.1. A 1-d problem with 2 parameters
$\begin{split}dx^{i}=\left(\gamma^{0}-x\right)dt+\left(0.5+\exp\left(-(x-\gamma^{1})^{2}\right)\right)dB,\quad\textnormal{where}\quad x_{0}\sim{\mathcal{N}}(0,1).\end{split}$
@@
The true parameter used to generate the data is $\gamma_{true}=[0,0]$. The initial parameter in the optimization is $\gamma_{0}=[5,1]$. The number of data points is $N_{data}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
@@
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The histograms of $y_{k}$’s and $x_{T,l}$’s for $\gamma_{0}$ and $\gamma_{10}$, the parameter after 10 updates, are shown in Figure˜3. As we can see, the two distributions get much closer during the optimization. More specifically, the error $\gamma-\gamma_{true}$ is reduced from $[5,1]$ to [0.05,0.09][-0.05,0.09]. The computational time for all 10 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
@@
Figure 3. Histograms of $y_{k}$’s and $x_{T,l}$’s. Left: plot at $\gamma_{0}$. Right: at $\gamma_{10}$.
5.2. A 5-d problem with 6 parameters
We try a 5-dimensional Lorenz 96 system. The SDE is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+\gamma^{i}-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left((x-\gamma^{M}\mathbbm{1})^{2}\right)\right)dB^{i},$
@@ 5.2. A 5-d problem with 6 parameters
The true parameter used to generate the data is $\gamma_{true}=[5,6,7,8,9,2]$. The initial parameter in the optimization is $\gamma_{0}=[0,0,0,0,0,0]$. The number of data points is $N_{data}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The history of the $l^{2}$-distance $|\gamma_{n}-\gamma_{true}|$ for $n=0$ to $50$ is shown in Figure˜4. As we can see, $\gamma_{n}$ gets closer to $\gamma_{true}$ during optimization, with $\gamma_{48}$ achieving the minimum distance. More specifically,
@@
The computational time for all 50 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
Figure 4. History of $|\gamma_{n}-\gamma_{true}|$ for the 5-d diffusion model.
@@ A.1. One-step results
The second term in Section˜A.1 can be very large for high-dimensional noise since it involves $M$. For now, we still use noise in all directions, so we try to avoid using the kernel formula for parameter derivative. The kernel formula for the spatial derivative does not have terms involving $M$ [24], so we can still use that. We will try to use low-dimensional noise in later papers, which will incur additional difficulties.
@@ A.3. Formal continuous-time limit
$\begin{split}\nabla\log h_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})+\int_{t=0}^{T}\frac{1}{\sigma_{t}(x_{t})}\delta F^{\gamma}(x_{n})\cdot dB\middle|x_{N}\right].\end{split}$

### v2

@@ 5.1. Problem and algorithm description
where the true parameters $\gamma_{true}:=[\gamma_{true}^{0},\ldots,\gamma_{true}^{N_{\gamma}-1}]$. The samples are $y:=x_{T}$ obtained at the end of independent paths. Denote the target distribution (the one we try to match) as $P$; when the target is empirical distribution prescribed by a set of data, $P$ is a sum of Dirac delta functions. Our framework can work on more general data not necessarily generated by an SDE, but here we restrict ourselves to this simple case, so that after optimization, we can check our result with the true parameters.
We want to fit several unknown parameters $\gamma=[\gamma^{0},\ldots,\gamma^{N_{\gamma}-1}]$ in the generative SDE model such that its marginal distribution matches the data. The prototype of the SDE is
@@
Denote the density at $T$ by $h^{\gamma}$ and measure by $H^{\gamma}$ – this is the distribution that we want to optimize.
In the current paper, we use simple models, such as one based on the Lorenz 96 system. These simple models can be interpreted as the result of the prior knowledge we had about how the data was generated. Such prior knowledge can be naturally incorporated into the SDE model. In the future, we will test on problems with less prior knowledge, so we will need more flexible models to express $F$ and $\sigma$, such as neural networks.
Also, note that the gradient used in the optimization is not computed by backpropagation; rather, it only involves a covector process that runs forward in time and uses much less memory. We also remark that we can choose to optimize the stationary distribution using Section˜3.4, which may be more efficient for some cases.
5.2. KL divergence as loss and gradient descent
Our training loss is the KL divergence, but we never need to evaluate the KL-divergence $D_{\mathrm{KL}}(P\|H^{\gamma})$; we only need its parameter-gradient. When $P$ admits a density $p$ (w.r.t. Lebesgue measure), the KL-divergence is
$D_{\mathrm{KL}}(P\|H^{\gamma})=\int p(y)\log\frac{p(y)}{h^{\gamma}(y)}\,dy=\underbrace{\int p(y)\log p(y)\,dy}_{\text{independent of }\gamma}-\int p(y)\log h^{\gamma}(y)\,dy.$
Hence, its parameter-derivative is
$\delta D_{\mathrm{KL}}(P\|H^{\gamma})=-\int_{\mathbb{R}^{M}}p(y)\delta\log h^{\gamma}(y)\,dy.$
If $P$ is given only through samples $\{y_{k}\}_{k=1}^{N_{\mathrm{data}}}\sim P$ (or is strictly singular), we optimize the same $\gamma$-dependent term (cross-entropy), namely
$\mathcal{L}(\gamma):=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\log h^{\gamma}(y_{k}),\qquad\delta\mathcal{L}(\gamma)=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\delta\log h^{\gamma}(y_{k}).$
We can compute the cross-entropy, but there is little benefit in doing so here, and it would require a few additional steps. In this paper, we judge the accuracy of the trained model by the discrepancy between the trained parameters and the ground-truth parameters; this criterion is complementary to the KL/cross-entropy.
Then we need to decide how to compute $\delta\log h^{\gamma}$ at the prescribed locations $y_{k}$. It is difficult to obtain paths ending exactly at $y_{k}$, so we need an approximation. For each $\gamma$ during optimization, we simulate $L$-many paths by this SDE, and we label each path by $l$, in particular, the terminal location is denoted by $x_{T,l}$. We approximate $y_{k}$ by its $N_{\mathrm{neighbor}}$-nearest neighbors in $\{x_{T,l}\}_{l=1}^{L}$, and therefore $\delta\log h^{\gamma}(y_{k})$ is approximated by the average of the divergence-kernel expression in Section˜1.1 over those $N_{\mathrm{neighbor}}$ paths. Hence,
(11) $\begin{split}\delta D_{\mathrm{KL}}(P\|H^{\gamma})\approx\hat{\delta}D_{\mathrm{KL}}(P\|H^{\gamma}):=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\frac{1}{N_{\mathrm{neighbor}}}\sum_{x_{T,l}\in N(y_{k})}\mathcal{T}(x_{T,l}),\end{split}$
where $N(y_{k})$ is the neighbors of $y_{k}$, and $\mathcal{T}(x_{T,l})$ is the expression in Section˜1.1 on the path ending at $x_{T,l}$.
After approximately computing the parameter-derivative, we update the parameter by gradient descent,
$\begin{split}\gamma_{n+1}=\gamma_{n}-\eta\hat{\delta}D_{\mathrm{KL}}(P\|H^{\gamma}).\end{split}$
@@ 5.2. KL divergence as loss and gradient descent
We can give a direct interpretation of the overall optimization strategy. Before each update, we simulate $L$ many sample paths, then we decide which of these paths’ terminal state is closest to the data (this is the effect of choosing $N_{\mathrm{neighbor}}$-nearest neighbors). On the other hand, the divergence-kernel linear response formula tells us how changing parameters affects the marginal log-densities at those terminal states. Combining both facts, we can now update parameters in the direction which increases the marginal log-densities on those terminal states close to the data. Hence, overall, each parameter update moves the marginal distribution of the SDE towards the data.
This interpretation also suggests that the best choice of $N_{\mathrm{neighbor}}$ is often $1$ or a small number. Indeed, increasing $N_{\mathrm{neighbor}}$ replaces the ideal conditional expectation $\mathbb{E}[\cdot\mid x_{T}=y_{k}]$ by a local average over nearby terminal states. While this can reduce Monte-Carlo variance, it also introduces a bias: points in high-density regions of the sample cloud tend to be selected as neighbors more frequently than points near the periphery. As a result, when $N_{\mathrm{neighbor}}$ is too large, the update may overweight central samples and encourage an artificial contraction of the learned distribution toward its center, even when the parameters are already close to optimal. For example, in the idealized situation where the simulated terminals match the data exactly, taking $N_{\mathrm{neighbor}}=1$ yields an unbiased 1-to-1 neighbor assignment, whereas $N_{\mathrm{neighbor}}>1$ mixes in off-target terminals and reintroduces geometric bias toward locally denser regions. Therefore we typically take $N_{\mathrm{neighbor}}=1$ (or a small value) unless additional smoothing is needed for stability.
We stress that the nearest-neighbor step is merely a simple numerical device for approximating the conditional expectation appearing in the divergence-kernel formula. One could replace it by other local averaging schemes without changing the underlying update principle. Therefore, the more substantive question is not the optimality of nearest-neighbor, but whether the resulting update direction admits the correct interpretation: namely, that it increases the model marginal log-density in regions supported by the data.
5.3. A 1D problem with 2 parameters
First, we try a 1-dimensional problem. The SDE model is
@@ 5.3. A 1D problem with 2 parameters
$\begin{split}dx=\left(\gamma^{0}-x\right)dt+\left(0.5+\exp\left(-(x-\gamma^{1})^{2}\right)\right)dB,\quad\textnormal{where}\quad x_{0}\sim{\mathcal{N}}(0,1).\end{split}$
@@
The true parameter used to generate the data is $\gamma_{true}=[0,0]$. The initial parameter in the optimization is $\gamma_{0}=[5,1]$. The number of data points is $N_{\mathrm{data}}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{\mathrm{neighbor}}=5$; we will test the effect of $N_{\mathrm{neighbor}}$ in the next example with higher dimension $M$.
Now we decide $\alpha$. For simplicity, we let it be a constant. To decide its value, recall that $\alpha$ is used to temper the growth of the covector process $\nu$. In view of Section˜1.1, the homogeneous part of $\nu$ grows like
$\begin{split}d\nu=(\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu dt+\cdots\end{split}$
If $\alpha$ is too small then $\nu$ would grow exponentially fast; we do not want this. For this 1D problem, we can directly compute these terms. We want the homogeneous part of $\nu$ to decay, which can be achieved if
$\begin{split}\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}=4\exp\left(-2(x-\gamma^{1})^{2}\right)(x-\gamma^{1})^{2}+1\leq 2/e+1\leq 3\leq\alpha\end{split}$
So we can just set $\alpha$ to be somewhat larger than $3$, for example, we set $\alpha=5$. For more complicated situations, we can just try a few $\alpha$ and pick the one whose corresponding $\nu$ is moderate.
@@
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The histograms of data $\{y_{k}\}$ and samples $\{x_{T,l}\}$ for $\gamma_{0}$ and $\gamma_{10}$, the parameter after 10 updates, are shown in Figure˜3. As we can see, the two distributions get much closer during the optimization. More specifically, the error $\gamma-\gamma_{true}$ is reduced from $[5,1]$ to [0.05,0.09][-0.05,0.09].
@@
Figure 3. Histograms of data $\{y_{k}\}$ and samples $\{x_{T,l}\}$. Left to right: plots at $\gamma_{0}$, $\gamma_{2}$, and $\gamma_{10}$.
5.4. A 5D problem with 6 parameters
We try a 5-dimensional Lorenz 96 system. The SDE model is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+\gamma^{i}-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left(-(x-\gamma^{M}\mathbbm{1})^{2}\right)\right)dB^{i},$
@@ 5.4. A 5D problem with 6 parameters
The true parameter used to generate the data is $\gamma_{true}=[5,6,7,8,9,2]$. The initial parameter in the optimization is $\gamma_{0}=[0,0,0,0,0,0]$. The number of data points is $N_{\mathrm{data}}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{\mathrm{neighbor}}=5$.
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The history of $|\gamma_{n}-\gamma_{true}|^{2}/N_{\gamma}$ for $n=0$ to $50$ is shown in Figure˜4. As we can see, $\gamma_{n}$ gets closer to $\gamma_{true}$ during optimization, with $\gamma_{48}$ achieving the minimum distance. More specifically,
@@
Figure 4. History of $|\gamma_{n}-\gamma_{true}|^{2}/N_{\gamma}$ for the 5D DK-SDE model.
5.5. A 20D problem with 21 parameters
We try a 20-dimensional system. More importantly, we run some convergence test against several hyper-parameters and record computer time. The SDE model is similar to the one in Section˜5.4, except for that now we set dimension $M=20$ and $N_{\gamma}=21$. We set $\gamma_{\mathrm{true}}\in\mathbb{R}^{N}_{\gamma}$ to ramp linearly from $5$ to $9$ over the first ten components and then ramp linearly back from $9$ to $5$ over the next ten components; additionally, we set the last entry $\gamma_{\mathrm{true},\,N_{\gamma}-1}=1$. The initial guess of all parameters are still zeros. The number of data points is $N_{\mathrm{data}}=200$.
We vary several hyperparameters; when sweeping one parameter, we keep the others fixed at their default values. In Figure˜5 we plot the convergence history of $|\gamma_{n}-\gamma_{\mathrm{true}}|^{2}/N_{\gamma}$ for $n=0,\dots,1000$ under different choices of hyperparameters. Specifically, we sweep $\eta$, $N_{\mathrm{neighbor}}$, and $L$, while holding the remaining two at the defaults $\eta=0.1$, $N_{\mathrm{neighbor}}=1$, and $L=3000$. Smaller $\eta$ leads to slower convergence but a better final error; in future work, it may therefore be beneficial to decrease $\eta$ according to a schedule. Using $N_{\mathrm{neighbor}}=1$ introduces some oscillations but yields the best overall result, which is encouraging since this choice does not depend on the dimension $M$. Finally, increasing the sample size $L$ improves the results, as expected.
Refer to captionRefer to captionRefer to caption
Figure 5. Convergence history for the 20D DK-SDE model. From left to right: sweep over learning rate $\eta$, $N_{\mathrm{neighbor}}$, and number of samples $L$.
Then we plot the data and generated samples under the default hyperparameters in Figure˜6. The wall time for 1000 updates on an RTX 5060 GPU is about 201 seconds. As intended, the sample distribution moves closer to the data distribution during optimization. More specifically, the scatter plots indicate good agreement at least in the first and second moments, which is a natural consequence of distributions getting closer.
Refer to captionRefer to captionRefer to caption
Figure 6. Scatter plot of $(x^{0},x^{1})$ for samples generated by the DK-SDE model and the given data. From left to right: samples generated by SDE with initial parameter, after 70 updates, after 500 updates.
Direct baselines are unavailable for this problem. Methods based on differentiating through trajectories (adjoint/tangent backpropagation) are numerically ineffective once the dynamics becomes unstable/nonhyperbolic (often already at moderate horizons in our setting), while likelihood-ratio constructions typically do not apply to parameters in the diffusion or suffer prohibitive variance. Other approaches (e.g. discrepancy or feature matching) optimize different objectives.
@@ A.1. One-step results
The second term in Section˜A.1 can be very large for high-dimensional noise since it involves $M$. For now, we still use noise in all directions, so we try to avoid using the kernel formula for parameter-derivative. The kernel formula for the spatial derivative does not have terms involving $M$ [31], so we can still use that. We will try to use low-dimensional noise in later papers, which will incur additional difficulties.
@@ A.3. Formal continuous-time limit
$\begin{split}\delta\log h_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})+\int_{t=0}^{T}\frac{1}{\sigma_{t}(x_{t})}\delta F^{\gamma}(x_{t})\cdot dB\middle|x_{T}\right].\end{split}$